OpenStudy (anonymous):

Is this already simplified ? sqrt 8 + sqrt 50 ?

OpenStudy (inkyvoyd):

no, completely factor 8 to get 2*2*2, and 50 to get 2*5*5. now see what you get.

OpenStudy (inkyvoyd):

Yes.

OpenStudy (inkyvoyd):

7sqrt 2 was the correct answer, but then you deleted it.

OpenStudy (anonymous):

\[\LARGE \sqrt{8}+\sqrt{50}=\sqrt{4\cdot 2}+\sqrt{25\cdot 2}=2\sqrt2 +5\sqrt2\]

OpenStudy (inkyvoyd):

=\[7\sqrt{2}\]

OpenStudy (anonymous):

@inkyvoyd you're right... \[\LARGE 7\sqrt2\] is the final answer :D